Your small dog gets away from you and begins running down the street. If your dog is running at a speed of 8.5 km/h and you can run 10 km/h, how many second' head start can you afford to give your dog if you want to catch him within 30 s of when you start after him?
Answer by Alan3354(69443)

Your small dog gets away from you and begins running down the street. If your dog is running at a speed of 8.5 km/h and you can run 10 km/h, how many second' head start can you afford to give your dog if you want to catch him within 30 s of when you start after him?
---------------------
Your speed relative to the dog is 1.5 km/h (10 - 8.5)
1.5 km = 1500 meters
1500 m/hr = 25 meters/min
30 sec = 0.5 minute
25 m/min for 0.5 min = 12.5 meters
12.5 meters head start
---------------------
12.5 meters @ 10 km/hr
10 km/hr = 10000/3600 m/sec = 25/9 m/sec
12.5 meters/(25/9 m/sec) = 4.5 seconds head start